# Large Language Models for Wireless Communications: From Adaptation to Autonomy

### v1

@@ V Challenges and Opportunities
Despite the great potential for the application of LLMs in communications, there are still many challenges that need to be addressed. This section identifies several research topics for further investigation.
@@ V-A Collaboration of LLM and Small Models
LLMs exhibit powerful generalization and reasoning capabilities but incur high computational cost and latency—making them ill-suited for real-time or resource-constrained tasks, particularly in the physical layer processing. A promising direction lies in designing collaborative frameworks where LLMs act as high-level coordinators while small domain-specific models handle fast, localized tasks, such as beam prediction, channel estimation, and power control. This hierarchical architecture can balance global reasoning with real-time responsiveness. Future research should also explore dynamic task delegation mechanisms, where the LLM selectively invokes or configures small models based on context and performance goals.
@@ V-B Multimodal Integration for Wireless Foundation Models
V-C Non-Linguistic Agentic Large AI Models
@@
V-D Toward Self-Improving and Lifelong Learning Agents
Wireless environments are inherently non-stationary: user mobility, time-varying interference patterns, and service demands change over time. While current agentic LLMs can respond to natural language prompts or feedback, they lack true self-improving capabilities, limiting long-term adaptability and robustness. Incorporating lifelong learning and memory consolidation into agentic wireless systems can enable continual adaptation to evolving network conditions. Techniques, such as meta-learning, episodic memory replay, and autonomous feedback generation, could support this functionality.

### v2

@@ V Challenges and Opportunities
Despite their significant promise, the application of LLMs in communications faces considerable challenges. This section outlines key directions for future research.
@@ V-A Collaboration of LLM and Small Models
LLMs exhibit powerful generalization and reasoning capabilities but incur high computational cost and latency—making them ill-suited for real-time or resource-constrained tasks, particularly in the physical layer processing. A promising direction lies in designing collaborative frameworks where LLMs act as high-level coordinators while small domain-specific models handle fast, localized tasks, such as beam prediction, channel estimation, and power control. This hierarchical architecture can balance global reasoning with real-time responsiveness. Future research should also explore dynamic task delegation mechanisms, where LLMs selectively invoke small models based on context and performance goals.
@@ V-B Multimodal Integration for Wireless Foundation Models
V-C Safety and Verification of Agentic LLMs
The generative and autonomous nature of agentic LLMs poses significant challenges in guaranteeing safe and predictable behavior in mission-critical networks. Existing reactive approaches, such as safety guardrails, can prevent basic violations on a per-action basis but are insufficient for ensuring the logical integrity of complex, multi-step plans. A promising direction lies in adopting formal verification, a proactive paradigm that can mathematically prove a plan’s adherence to critical system invariants before execution. Future research should develop scalable verification algorithms for long-horizon tasks and robust methods for translating ambiguous, LLM-generated plans into verifiable formal models.
V-D Non-Linguistic Agentic Large AI Models